# Conformal prediction with local weights: randomization enables robust guarantees

### v3

@@ 5.1.2 Multivariate setting
5.2 Real data example: predicting medical insurance charges
@@
Now, we will compare the performance of the local CP methods on a real data. We work with a medical insurance dataset [^5] The data from this experiment was obtained from https://www.kaggle.com/datasets/rajgupta2019/medical-insurance-dataset. with $3630$ data points, where the response charges is the cost of the medical charges for that patient, and the features consist of the patient’s age, sex, BMI, smoking status (smoker or nonsmoker), region of residence within the U.S. (e.g., Northeast), and number of children. For our experiment, we will use $Y=\log(\texttt{charges})$ as the response, and $X_{\textnormal{age}}$, $X_{\textnormal{sex}}$, $X_{\textnormal{BMI}}$, $X_{\textnormal{smoking}}$ as the features. After splitting the data equally into a pretraining set, a calibration set, and a test set, [^6] Here we depart from the terminology of the earlier sections in our paper—the term calibration set refers to the data $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ on which we implement the methods baseLCP, calLCP, and RLCP, using a predefined score function $s$ that is based on a pretrained predictor $\widehat{f}$. This same set was previously called the training set throughout the paper, but here we refer to it as the calibration set to distinguish it from the pretraining set used for fitting $\widehat{f}$. we train three different base predictors $\widehat{f}$ on the pretraining set, using a linear regression, a $2$-layer neural network, or a random forest. For each $\widehat{f}$ we define the corresponding pretrained score function $s(x,y)=|y-\widehat{f}(x)|$.
Next, we will define the localization kernel $H$ to place high weight for pairs of patients that have the same sex and smoking status, and similar values for age and BMI: fixing a choice of bandwidths $h_{\texttt{age}}$ and $h_{\texttt{BMI}}$, we define
$H(x,x^{\prime})=\frac{1}{4h_{\texttt{age}}h_{\texttt{BMI}}}\cdot\mathbbm{1}\{x_{\texttt{sex}}=x^{\prime}_{\texttt{sex}},x_{\texttt{smoking}}=x^{\prime}_{\texttt{smoking}},|x_{\texttt{age}}-x^{\prime}_{\texttt{age}}|\leq h_{\texttt{age}},|x_{\texttt{BMI}}-x^{\prime}_{\texttt{BMI}}|\leq h_{\texttt{BMI}}\}.$
This satisfies the condition (12), i.e., that $H(x,\cdot)$ defines a density with respect to an appropriate base measure: namely, sampling from $H(x,\cdot)$ consists of sampling an age uniformly from $[x_{\textnormal{age}}-h_{\textnormal{age}},x_{\textnormal{age}}+h_{\textnormal{age}}]$, sampling a value of BMI uniformly from $[x_{\textnormal{BMI}}-h_{\textnormal{BMI}},x_{\textnormal{BMI}}+h_{\textnormal{BMI}}]$, and copying the values of $x_{\textnormal{sex}}$ and $x_{\textnormal{smoking}}$. In our implementation, for simplicity we will take $h_{\textnormal{age}}=h_{\textnormal{BMI}}$ to always equal the same value, chosen from $h\in\{2,3,\dots,7,8\}.$
@@
Figure 5: Marginal coverage (averaged over the test set) and prediction interval width (median over the test set), for each of the three methods across different bandwidths and for different choices of the pretrained base predictor $\widehat{f}$, for the real data experiment. Results are averaged over $100$ random splits of the data into pretraining, calibration, and test sets. See Section 5.2 for details.
@@
In Figure 5, for each of the three localized CP methods, we present the marginal coverage and the median width of the prediction intervals on the test set, averaged over $100$ random splits of the data. While RLCP and calLCP attains exact target level of marginal coverage, baseLCP prediction intervals are conservative for our candidate bandwidth choices. We can see that RLCP produces wider prediction intervals than calLCP at the lower values of $h$, while producing similar or slightly narrower intervals for higher $h$, for the various base algorithm choices. We will see next that RLCP offers better empirical conditional coverage at lower bandwidth values, which may justify the need for its larger width at these low values of $h$.
@@
In this real data setting, there are many practical questions we can ask regarding test-conditional coverage. In particular, since each data point is a patient, test-conditional coverage can be interpreted as asking whether the predictive intervals are equally valid across different subgroups of patients (e.g., younger vs older). Here we aim to gain some insight into these questions by examining the empirical coverage of each method conditional on certain patient characteristics. Ideally, we would want a method to show approximately constant conditional coverage across different subgroups of patients. In the left panel of Figure 6, we show coverage conditional on smoking status, i.e., conditional on the value of the categorical covariate $X_{\texttt{smoking}}$. The bar plot shows the value of ${\mathbb{P}}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\mid(X_{n+1})_{\texttt{smoking}}=x\}$ for each $x\in\{\texttt{yes},\texttt{no}\}$. While RLCP shows approximately equal coverage levels for the two groups across all three base predictors $\widehat{f}$, for calLCP the performance is quite variable, with extremely disparate coverage levels for some choices of the base algorithm. In the right panel of Figure 6, we ask the same question with respect to $X_{\texttt{BMI}}$. Since this is a continuous covariate, we cannot compute conditional coverage exactly, but instead estimate it with a sliding window whose width is chosen to have $5\%$ mass: that is, we are plotting ${\mathbb{P}}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\mid(X_{n+1})_{\texttt{BMI}}\in[x-\Delta_{0},x+\Delta_{1})\}$ where, for each $x$, $\Delta_{0}$ and $\Delta_{1}$ are chosen so that $2.5\%$ of the population has BMI in the range $[x-\Delta_{0},x)$, and same for the range $[x,x+\Delta_{1})$. Here we again see that RLCP shows better conditional coverage compared to calLCP particularly at lower bandwidth $h=2$—the estimated conditional coverage for RLCP lies closer to the target level of $90\%$ while for calLCP we see slightly more prominent over- or undercoverage at various values of BMI.
Figure 6: Comparison of local coverage of the three methods, conditioning on $X_{\texttt{smoking}}$ (left) and on $X_{\texttt{BMI}}$ (right). See Section 5.2 for details.

### v4

@@ 5.1.2 Multivariate setting
5.2 Real data example: predicting age of abalones
Next, we will compare the performance of the local CP methods on a real dataset on abalones (a species of mollusk) (Nash et al.,, 1994). [^5] The data for this experiment was obtained from https://archive.ics.uci.edu/dataset/1/abalone. The objective is to predict the age of abalones based on their physical measurements. Since the economic value of abalone is positively correlated with its age, but determining age precisely requires a costly procedure, accurately predicting age from physical characteristics alone becomes both a valuable and practical problem. This dataset has $4177$ data points, where the response rings determines the age of the abalone, and the features consist of the abalone’s sex (categorized as Male, Female, or Infant), length (the longest shell measurement), diameter (perpendicular to length), height, whole weight, shucked weight (i.e., weight of the meat inside), viscera weight (i.e., gut weight after bleeding) and shell weight (the weight after drying).
@@ 5.2 Real data example: predicting age of abalones
For our experiment, we will use $Y=\texttt{ring}$ as the response, and $X_{\texttt{sex}}$, $X_{\texttt{length}}$, $X_{\texttt{diameter}}$, $X_{\texttt{height}}$ and $X_{\texttt{whole\_weight}}$ as the features. After splitting the data equally into a pretraining set, a calibration set, and a test set, [^6] Here we depart from the terminology of the earlier sections in our paper—the term calibration set refers to the data $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ on which we implement the methods baseLCP, calLCP, and RLCP, using a predefined score function $s$ that is based on a pretrained predictor $\widehat{f}$. This same set was previously called the training set throughout the paper, but here we refer to it as the calibration set to distinguish it from the pretraining set used for fitting $\widehat{f}$. we train three different base predictors $\widehat{f}$ on the pretraining set, using a linear regression, a $2$-layer neural network, or a random forest. For each $\widehat{f}$ we define the corresponding pretrained score function $s(x,y)=|y-\widehat{f}(x)|$. Next, we will define the localization kernel $H$ to place high weight for pairs of patients that have the same sex, and similar values for length, diameter, height and whole weight: fixing a choice of bandwidths $h_{\texttt{length}}$, $h_{\texttt{diameter}}$, $h_{\texttt{height}}$ and $h_{\texttt{whole\_weight}}$, we define
$H(x,x^{\prime})=\frac{1}{2h_{\texttt{sex}}\cdot 2h_{\texttt{diameter}}\cdot 2h_{\texttt{height}}\cdot 2h_{\texttt{whole\_weight}}}\cdot\mathbbm{1}\bigg{\{}x_{\texttt{sex}}=x^{\prime}_{\texttt{sex}},|x_{\texttt{length}}-x^{\prime}_{\texttt{length}}|\leq h_{\texttt{length}}\\ \hskip 142.26378pt|x_{\texttt{diameter}}-x^{\prime}_{\texttt{diameter}}|\leq h_{\texttt{diameter}},|x_{\texttt{height}}-x^{\prime}_{\texttt{height}}|\leq h_{\texttt{height}},\\ |x_{\texttt{whole\_weight}}-x^{\prime}_{\texttt{whole\_weight}}|\leq h_{\texttt{whole\_weight}}\bigg{\}}.$
[…]
@@
Figure 5: Marginal coverage (averaged over the test set) and prediction interval width (median over the test set), for each of the three methods across different bandwidths and for different choices of the pretrained base predictor $\widehat{f}$, for the real data experiment. Results are averaged over $50$ random splits of the data into pretraining, calibration, and test sets. See Section 5.2 for details.
@@
In Figure 5, for each of the three localized CP methods, we present the marginal coverage and the median width of the prediction intervals on the test set, averaged over $50$ random splits of the data. While RLCP and calLCP attains exact target level of marginal coverage, coverage of baseLCP prediction intervals are lower and higher respectively for smaller and larger values of $h$. We further see that median length of prediction intervals are slightly narrower for RLCP prediction intervals for most bandwidth choices and for the various base algorithm choices. We will see next that RLCP offers better empirical conditional coverage at lower bandwidth values, which may justify the need for the prediction intervals with slightly larger width at the lowest values of $h$.
@@
In this real data setting, there are many practical questions we can ask regarding test-conditional coverage. In particular, since each data point corresponds to an individual abalone, test-conditional coverage can be interpreted as asking whether the predictive intervals are equally valid across different subgroups of abalones (e.g., smaller vs larger abalones). Here we aim to gain some insight into these questions by examining the empirical coverage of each method conditional on certain physical measurement for abalone. Ideally, we would expect approximately constant conditional coverage across different subgroups of abalones. In the left panel of Figure 6, we show coverage conditional on sex, i.e., conditional on the value of the three levels of covariate $X_{\texttt{sex}}$. The bar plot shows the value of ${\mathbb{P}}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\mid(X_{n+1})_{\texttt{sex}}=x\}$ for each $x\in\{\texttt{Male},\texttt{Female},\texttt{Infant}\}$. While RLCP shows approximately equal coverage levels for the three groups across all three base predictors $\widehat{f}$, for calLCP the performance is quite variable, with extremely disparate coverage levels for some choices of the base algorithm.
In the right panel of Figure 6, we ask the same question with respect to $X_{\texttt{length}}$. Since this is a continuous covariate, we cannot compute conditional coverage exactly, but instead estimate it with a sliding window whose width is chosen to have $5\%$ mass: that is, we are plotting ${\mathbb{P}}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\mid(X_{n+1})_{\texttt{length}}\in[x-\Delta_{0},x+\Delta_{1})\}$ where, for each $x$, $\Delta_{0}$ and $\Delta_{1}$ are chosen so that $2.5\%$ of the abalone population has length in the range $[x-\Delta_{0},x)$, and same for the range $[x,x+\Delta_{1})$. Here we again see that RLCP shows better conditional coverage compared to calLCP particularly at lower bandwidth $h=0.05$—the estimated conditional coverage for RLCP is approximately constant at the target level of $90\%$ while for calLCP, we see very prominent over- or undercoverage at various values of length.
Figure 6: Comparison of local coverage of the three methods, conditioning on $X_{\texttt{sex}}$ (left) and on $X_{\texttt{length}}$ (right). See Section 5.2 for details.